# Stable subgroups of graph products

### v1

@@ Abstract
We extend the characterization of stable subgroups of right-angled Artin groups from [RAAGstable] to the case of graph products of infinite groups. Specifically, we show that the stable subgroups of such graph products are exactly the subgroups that quasi-isometrically embed in the associated contact graph. Equivalently, they are the subgroups that satisfy a condition arising from the defining graph: a stable subgroup is an almost join-free subgroup. In particular, we recover the equivalence between stable and purely loxodromic subgroups of [RAAGstable] in the case where all torsion subgroups of the vertex groups are finite.
@@ 1 Introduction
Given a simplicial graph $\Gamma$ and a family of finitely generated groups $\mathcal{G}=\{G_{v}\mid v\in V(\Gamma)\}$ indexed by the vertices of $\Gamma$, the associated graph product $\Gamma\mathcal{G}$ is generated by the vertex groups $G_{v}$, with the additional relation that elements of adjacent vertex groups commute (as detailed in Section 2.3). Graph products were introduced by Green in her PhD thesis [greennormalform] and have since been widely studied. They provide an interplay between free products and direct products of groups, and generalize both right-angled Artin groups (RAAGs) and right-angled Coxeter groups (RACGs). In this paper, we focus on the case where the defining graph $\Gamma$ has no isolated vertices, and where the groups $G_{v}$ are all finitely generated and infinite; this situation generalizes the one of RAAGs.
@@
This notion of stability was introduced in [DT] to generalize both the quasi-convex subgroups of hyperbolic groups, and the convex cocompact subgroups of mapping class groups [FarbMosher]. They have since been well studied for a broad collection of groups, see for example [sistoHypEmbed, CordesDurham, AMST, HKim, abbott2022acylindrically]. In particular, [RAAGstable] proved a geometric characterization of the stable subgroups of RAAGs.
@@
[RAAGstable, Theo 1.1] Let $\Gamma$ be a finite connected graph and $H$ a finitely generated subgroup of the right angled Artin group $A(\Gamma)$. The following are equivalent:
@@
In this theorem stability is characterized in terms of action of the group on the extension graph. This space was introduced by Kim and Koberda in [KimKoberda], and is a hyperbolic graph on which $A(\Gamma)$ acts by isometries [KimKoberda2]. In the language of [BCKMT], the equivalence of (1) and (2) in 1.1 implies that the extension graph is a universal recognizing space for the stable subgroups of $A(\Gamma)$. A universal recognizing space for a group $G$ is a hyperbolic space on which $G$ acts by isometries, and in which all the stable subgroups embed quasi-isometrically through their orbit maps for this action. This has also been called a stability recognizing space, with the additional (reverse) assumption that every subgroup quasi-isometrically embedded through its orbit maps is stable [CDEK].
A universal recognizing space for a group can be expected to play a role analogous to the one of the curve complex for mapping class groups. In fact, 1.1 is also true for mapping class groups when the extension graph is replaced by the curve complex and with the additional assumption in condition (3) that the group $H$ is undistorted [MR2465691, Hamenstaedt2005WordHE, BBKL]. Conditions (1) and (2) are also more generally equivalent for hierarchically hyperbolic groups acting on the hyperbolic space associated to the maximal domain of some specific structure on the group [ABD], as well as for CAT(0) groups acting on the curtain model [CAT0], Morse-dichotomous groups acting on the contraction space [Zbinden, PetytZalloum], and certain relatively hyperbolic groups acting on the coned-off Cayley graph [ADT]. For graph products the analog of the curve graph that we are interested in is the contact graph of the prism-complex; one can refer to [Oyakawa] or [genevois2017] for other possibilities.
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices. Let $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of finitely generated infinite groups, and $P(\Gamma\mathcal{G})$ the prism complex of the graph product $\Gamma\mathcal{G}$. Let $H$ be a finitely generated subgroup of the graph product $\Gamma\mathcal{G}$. Then the following are equivalent.
@@
The contact graph $\mathcal{C}P(\Gamma\mathcal{G})$ is known to be a quasi-tree, and hence a hyperbolic space [valiunas2021graphproducts, Corollary C]. It is therefore a universal recognizing space for $\Gamma\mathcal{G}$ under the conditions of 1.2. It was also shown in [valiunas2021graphproducts, Corollary D] that, under these conditions, the action of $\Gamma\mathcal{G}$ on the contact graph is the largest acylindrical action. On the other hand, when isolated vertices are allowed, the authors showed in [BCKMT] that the largest acylindrical action of such a graph product does not necessarily provide a universal recognizing space for stable subgroups.
We note that the final equivalent condition in 1.2 is different from the one in Theorem 1.1. A subgroup of a graph product is called a join subgroup if it decomposes as a direct product of subgroups generated by disjoint subgraphs of the defining graph. The stable subgroups are then exactly the almost join-free subgroups, in other words those admitting only finite intersections with conjugates of join subgroups. (Note that our definition of (almost) join-free subgroups differs slightly from the one given by Tran in [Tran]). In the case of RAAGs, the purely loxodromic condition is known to be equivalent to asking $H$ to be (almost) join-free [KimKoberda2], and this equivalence is a key part of the proof of 1.1.
The reason this characterization cannot be extended to general graph products is the potential existence of infinite torsion subgroups. Under the added assumption that such subgroups do not exist, we are able to recover the final equivalent statement of 1.1; also, in this specific case our notion of almost join-free subgroups corresponds with the one of Tran [Tran]. In the case where infinite torsion subgroups are allowed, stable implies purely loxodromic, but not the other way around (see Section 5).
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices. Let $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of infinite finitely generated groups with no infinite torsion subgroups, and $P(\Gamma\mathcal{G})$ the prism complex of the graph product $\Gamma\mathcal{G}$. Let $H$ be a finitely generated subgroup of the graph product $\Gamma\mathcal{G}$. Then the following are equivalent.
  1. 1.
    $H$ is stable in $\Gamma\mathcal{G}$.
  2. 2.
  3. 3.
    $H$ is almost join-free.
  4. 4.
    $H$ is purely loxodromic with respect to the action on $\mathcal{C}P(\Gamma\mathcal{G})$.
We remark that the hypotheses on the vertex groups and on the graph in 1.2 and 1.3 are necessary. Firstly, if $\Gamma$ contains an isolated vertex then the associated vertex group will always be almost join-free, while its orbit in the contact graph $\mathcal{C}P(\Gamma\mathcal{G})$ will always have finite diameter, and the vertex group will be stable in $\Gamma\mathcal{G}$ if and only if it is hyperbolic. Secondly, if finite vertex groups were allowed, we could consider the case where $\Gamma$ is a path with two vertices, and the vertex groups are $\mathbb{Z}$ and $\mathbb{Z}_{2}$. This would mean that that $\Gamma\mathcal{G}$ is an infinite join group (so not almost join-free), while being stable within itself, and with $\mathcal{C}P(\Gamma\mathcal{G})$ having finite diameter.
Note that this hypothesis on vertex groups excludes RACGs from our theorem; in that case an analogous version of the curve complex is given by the the graph of maximal product regions introduced by Oh [Oh], and under some hypotheses Cashen, Dani, Edletzberger, and Karrer prove the equivalence between (1) and (2) for this graph [CDEK, Cor 6.10]. As RACGs are hierarchically hyperbolic groups, the work in [ABD] also applies. It is therefore natural to ask if a more general result exists for graph products.
Question 1.4.
@@
In [RAAGstable], as a consequence of the characterizations of stable subgroups in RAAGs, it is also shown that stable subgroups of RAAGs are free [RAAGstable, Prop 8.1, 8.2]. In the case of graph products, we get the following generalization as a direct consequence of 1.2, the contact graph being a quasi-tree, and the fact that groups which quasi-isometrically embed into quasi-trees are virtually free (see for example [Button2023, Corollary 10.7]).
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices and $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ a collection of finitely generated infinite groups. The stable subgroups of $\Gamma\mathcal{G}$ are virtually free.
In the non-hyperbolic cases of both RAAGs of connected graphs and mapping class groups, there is an additional characterization of stable subgroups: they are exactly the infinite-index Morse subgroups [HKim, Tran2019]. This is not the case for RACGs [Tran2019, Theorem 1.11 and Corollary 1.12], however it could plausibly hold for graph products of infinite groups.
Question 1.6.
Let $\Gamma$ be a finite connected simple graph and $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ a collection of finitely generated infinite groups. Do the stable subgroups of $\Gamma\mathcal{G}$ coincide with the infinite-index Morse subgroups?
Remark 1.7.
In private correspondence, Anthony Genevois suggested an alternate proof of 1.5 using Stallings-Dunwoody decompositions that would apply more generally to non-torsion infinite-index accessible Morse subgroups, further motivating 1.6.
@@
Acknowledgments. The authors are grateful to the organizers of the 2023 Women in Groups, Geometry, and Dynamics program, where our collaboration started. We would like to thank Anthony Genevois for his feedback, both at the beginning of this project and on a later draft, and for his enlightening comments on [Genevois2019]. Our thanks also to Chris Cashen and Annette Karrer for their explanation of the case of right-angled Coxeter groups.
@@ 2.2 Stable subgroups
  • For each choice of constants $\lambda\geq 1,c\geq 0$, there is a constant $M>0$ (depending on $\lambda,c$) such that any two $(\lambda,c)$–quasi-geodesics in $G$ with the same end points in $H$ are each contained in the MM–neighborhood of the other.
Note that stability is independent of the choice of finite generating sets for $H$ and $G$ [DT]. In particular, it is preserved under group automorphisms such as conjugation. We note this explicitly because much of the work in this paper is dedicated to the study of condition (3) in 1.2 which deals with conjugacy. Namely, we have:
@@ 2.4 Associated graphs and complexes
Throughout this paper we consider several different Cayley graphs associated to $\Gamma\mathcal{G}$ and their corresponding metrics. For more details on these the reader can refer to [genevois2017] or [valiunas2021graphproducts]. We recall here the basic definitions and facts we will need.
@@ 2.4.1 Prism Cayley graph and complex
Most of the work in this document uses what we call the prism Cayley graph of $\Gamma\mathcal{G}$, following Genevois’ study of graph products and other groups in [genevois2017]. This is the Cayley graph using the generating set $S_{p}$ (the prism generating set for $\Gamma\mathcal{G}$) defined as follows:
@@
The prism Cayley graph is denoted $\text{Cay}_{p}(\Gamma\mathcal{G})$, geodesic words for the prism generating set are referred to as prism geodesics or $S_{p}$–geodesics, and $|\cdot|_{p}$ is the prism length for a group element. Genevois proved that the prism Cayley graph of a graph product is always a quasi-median graph [genevois2017, Proposition 8.2].
@@
In this paper we refer to such a cell complex as the prism complex of $\Gamma\mathcal{G}$ for convenience, however they are in fact a subclass of what Genevois calls prism complexes of quasi-median graphs [genevois2017]. In fact, what we call the prism complex is the 2–skeleton of Genevois’ quasi-median complex, which is obtained by filling every clique with a simplex, and filling every 1–skeleton of an nn–cube with an nn–cube. As we do not need to consider any higher dimensional cells, we omit them here.
@@ 2.4.2 Hyperplanes and carriers
A hyperplane may also be identified with its carrier: the union of closed cells that intersect the geometric hyperplane. From definitions one can verify that the 0–skeleton of a carrier is precisely some coset of $G_{\mathsf{Star}(v)}$, where $v$ is the vertex associated to the hyperplane [genevois2017, GenevoisMartin]. We also make use of its 1–skeleton, the full subgraph induced in $P(\Gamma\mathcal{G})$ by these vertices (following [valiunas2021graphproducts]). Note that the edges of the 1–skeleton that do not belong to the hyperplane itself are labeled by generators from $G_{\mathsf{Link}(v)}$.
@@ 2.4.3 Contact graph
The definition above is parallel to the construction of the contact graph of a CAT(0) cube complex, introduced by Hagen in [Hagen]. Several results from the cube complex case also hold for graph products, in particular the following is known.
@@
[valiunas2021graphproducts, Corollary C] Let $\Gamma$ be a bounded degree simplicial graph, and let $\Gamma\mathcal{G}$ be a graph product of non-trivial groups. Then $\mathcal{C}P(\Gamma\mathcal{G})$ is a quasi-tree (and therefore a hyperbolic space), and the action of $\Gamma\mathcal{G}$ on $\mathcal{C}P(\Gamma\mathcal{G})$ is acylindrical.

### v2

@@ Abstract
We extend the characterization of stable subgroups of right-angled Artin groups from [28] to the case of graph products of infinite groups. Specifically, we show that the stable subgroups of such graph products are exactly the subgroups that quasi-isometrically embed in the associated contact graph. Equivalently, they are the subgroups that satisfy a condition arising from the defining graph: a stable subgroup is an almost join-free subgroup. In particular, we generalize the equivalence between stable and purely loxodromic subgroups from [28] in the case where all torsion subgroups of the vertex groups are finite, and the equivalence between stable and infinite index Morse subgroups from [36] in the case where the defining graph is connected.
@@ 1 Introduction
Given a simplicial graph $\Gamma$ and a family of finitely generated groups $\mathcal{G}=\{G_{v}\mid v\in V(\Gamma)\}$ indexed by the vertices of $\Gamma$, the associated graph product $\Gamma\mathcal{G}$ is generated by the vertex groups $G_{v}$, with the additional relation that elements of adjacent vertex groups commute (as detailed in Section 2.3). Graph products were introduced by Green in her PhD thesis [20] and have since been widely studied. They provide an interplay between free products and direct products of groups, and generalize both right-angled Artin groups (RAAGs) and right-angled Coxeter groups (RACGs). In this paper, we focus on the case where the defining graph $\Gamma$ has no isolated vertices, and where the groups $G_{v}$ are all finitely generated and infinite; this situation generalizes the one of RAAGs.
@@
This notion of stability was introduced in [12] to generalize both the quasi-convex subgroups of hyperbolic groups, and the convex cocompact subgroups of mapping class groups [13]. They have since been well studied for a broad collection of groups; see for example [35, 10, 3, 24, 2]. In particular, [28] proved a geometric characterization of the stable subgroups of RAAGs.
@@
[28, Theo 1.1] Let $\Gamma$ be a finite connected graph and $H$ a finitely generated subgroup of the right angled Artin group $A(\Gamma)$. The following are equivalent:
@@
In this theorem stability is characterized by the action of the subgroup on the extension graph of $A(\Gamma)$. This space was introduced by Kim and Koberda in [25], and is a hyperbolic graph on which $A(\Gamma)$ acts by isometries [26]. In the language of [5], the equivalence of (1) and (2) in 1.1 implies that the extension graph is a universal recognizing space for the stable subgroups of $A(\Gamma)$. A universal recognizing space for a group $G$ is a hyperbolic space on which $G$ acts by isometries, and in which all the stable subgroups embed quasi-isometrically through their orbit maps for this action. With the additional (converse) assumption that every subgroup quasi-isometrically embedded through its orbit maps is stable, this has also been called a stability recognizing space [9].
A universal recognizing space for a group can be expected to play a role analogous to the one of the curve complex for mapping class groups. In fact, 1.1 is also true for mapping class groups when the extension graph is replaced by the curve complex and with the additional assumption in condition (3) that the group $H$ is undistorted [23, 22, 6]. More generally, conditions (1) and (2) are equivalent for hierarchically hyperbolic groups acting on the hyperbolic space associated to the maximal domain of some specific structure on the group [1], as well as for CAT(0) groups acting on the curtain model [33], Morse-dichotomous groups acting on the contraction space [39, 34], and certain relatively hyperbolic groups acting on the coned-off Cayley graph [4].
For graph products the analog of the curve graph that we are interested in is the contact graph of the prism-complex; one can refer to [31] or [17] for other possibilities. Our main theorem is as follows.
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices, $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of finitely generated infinite groups, and $P(\Gamma\mathcal{G})$ be the prism complex of the graph product $\Gamma\mathcal{G}$. Let $H$ be a finitely generated subgroup of the graph product $\Gamma\mathcal{G}$. Then the following are equivalent.
@@
The contact graph $\mathcal{C}P(\Gamma\mathcal{G})$ is known to be a quasi-tree, and hence a hyperbolic space [38, Corollary C]. It is therefore a universal recognizing space for $\Gamma\mathcal{G}$ under the conditions of 1.2.
Remark 1.3.
The equivalence of (1) and (2) in 1.2 has also been proved independently (and simultaneously) by Joshua Perlmutter using different methods [32, Corollary 5.4 and Proposition 5.5], namely via a broader result for relatively hierarchically hyperbolic spaces [32, Theorem 1.3].
We can relate the final equivalent condition in 1.2 to the last condition in Theorem 1.1. A subgroup of a graph product is called a join subgroup if it decomposes as a direct product of subgroups generated by disjoint subgraphs of the defining graph. 1.2 says the stable subgroups are exactly the almost join-free subgroups, meaning those admitting only finite intersections with conjugates of join subgroups. (Note, our definition of almost join-free subgroups differs slightly from the one given by Tran in [37]). In the case of RAAGs, the purely loxodromic condition is known to be equivalent to asking $H$ to be (almost) join-free [26], and this equivalence is a key part of the proof of 1.1. This characterization cannot extend to general graph products due to the potential existence of infinite torsion subgroups, which are automatically purely loxodromic, but never stable (Section 5 details why stable implies purely loxodromic, but not the other way around). Under the added assumption that such subgroups do not exist, we recover the final equivalent statement of 1.1.
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices, $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of infinite finitely generated groups with no infinite torsion subgroups, and $P(\Gamma\mathcal{G})$ be the prism complex of the graph product $\Gamma\mathcal{G}$. A finitely generated subgroup $H$ of $\Gamma\mathcal{G}$ is stable if and only if it is purely loxodromic with respect to the action on the contact graph $\mathcal{C}P(\Gamma\mathcal{G})$.
In the non-hyperbolic cases of both $A(\Gamma)$ for connected $\Gamma$ and mapping class groups, there is another characterization of stable subgroups: they are exactly the infinite-index Morse subgroups [36, 24]. This is not the case for RACGs [36, Theorem 1.11 and Corollary 1.12], however it does hold for graph products of infinite groups, when the defining graph is connected.
Theorem 1.5.
Let $\Gamma$ be a finite connected simple graph with at least two vertices, and $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of finitely generated infinite groups. A subgroup $H$ of the graph product $\Gamma\mathcal{G}$ is stable if and only if it is infinite index and Morse.
The hypotheses on the vertex groups and on the graphs in the above theorems are necessary. For 1.5, if $\Gamma$ is not connected, then $\Gamma\mathcal{G}$ is a free product of infinite groups, with each of these free factors being both infinite index and Morse. At the same time, those free factors will be stable if and only if they are hyperbolic. 1.2 and 1.4 assume a slightly weaker hypothesis, namely that there are no isolated vertices in $\Gamma$. If $\Gamma$ contains an isolated vertex, then the associated vertex group will always be almost join-free, while its orbit in the contact graph will have finite diameter. This vertex group is also a free factor, and therefore Morse, so as before it will be stable if and only if it is hyperbolic. Moreover, it was shown in [38, Corollary D] that, under the conditions of 1.2, the action of $\Gamma\mathcal{G}$ on the contact graph is the largest acylindrical action. When isolated vertices are allowed, the authors showed in [5] that the largest acylindrical action of such a graph product does not necessarily provide a universal recognizing space for stable subgroups.
The other important hypothesis in the above theorems is that the vertex groups are infinite. Otherwise, we could consider the case where $\Gamma$ is a path with two vertices, and the vertex groups are $\mathbb{Z}$ and $\mathbb{Z}_{2}$. Here $\Gamma\mathcal{G}$ is an infinite join group (so not almost join-free), while being stable and finite index within itself, and with $\mathcal{C}P(\Gamma\mathcal{G})$ having finite diameter. Note that this hypothesis on vertex groups excludes RACGs from our theorem; in that case an analogous version of the curve complex is given by the the graph of maximal product regions introduced by Oh [30], and under some hypotheses Cashen, Dani, Edletzberger, and Karrer prove the equivalence between (1) and (2) for this graph [9, Cor 6.10]. As RACGs are hierarchically hyperbolic groups, the work in [1] also applies. It is therefore natural to ask if a more general result exists for graph products.
Question 1.6.
@@
Our final observation concerns what stability implies about subgroup type. Stable subgroups of RAAGs are free, by 1.1 and [28, Prop 8.1, 8.2]. For graph products, we get the following generalization as a direct consequence of 1.2, the contact graph being a quasi-tree, and the fact that groups which quasi-isometrically embed into quasi-trees are virtually free (see for example [8, Corollary 10.7]).
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices, and $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of finitely generated infinite groups. The stable subgroups of the graph product $\Gamma\mathcal{G}$ are virtually free.
@@
Acknowledgments. The authors are grateful to the organizers of the 2023 Women in Groups, Geometry, and Dynamics program, where our collaboration started. We would like to thank Anthony Genevois for his feedback, both at the beginning of this project and on a later draft, and for his enlightening comments on [18]. Our thanks also to Chris Cashen and Annette Karrer for their explanation of the case of right-angled Coxeter groups.
@@ 2.2 Stable subgroups
  • For each choice of constants $\lambda\geq 1,c\geq 0$, there is a constant $M\geqslant 0$ (depending on $\lambda,c$) such that any two $(\lambda,c)$–quasi-geodesics in $G$ with the same end points in $H$ are each contained in the MM–neighborhood of the other.
Note that stability is independent of the choice of finite generating sets for $H$ and $G$ [12]. In particular, it is preserved under group automorphisms such as conjugation. We note this explicitly because much of the work in this paper is dedicated to the study of condition (3) in 1.2 which deals with conjugacy. Namely, we have:
@@ 2.4 Associated graphs and complexes
Throughout this paper we consider several different Cayley graphs associated to $\Gamma\mathcal{G}$ and their corresponding metrics. For more details on these the reader can refer to [17] or [38]. We recall here the basic definitions and facts we will need.
@@ 2.4.1 Prism Cayley graph and complex
Most of the work in this document uses what we call the prism Cayley graph of $\Gamma\mathcal{G}$, following Genevois’ study of graph products and other groups in [17]. This is the Cayley graph using the generating set $S_{p}$ (the prism generating set for $\Gamma\mathcal{G}$) defined as follows:
@@
The prism Cayley graph is denoted $\text{Cay}_{p}(\Gamma\mathcal{G})$, geodesic words for the prism generating set are referred to as prism geodesics or $S_{p}$–geodesics, and $|\cdot|_{p}$ is the prism length for a group element. Genevois proved that the prism Cayley graph of a graph product is always a quasi-median graph [17, Proposition 8.2].
@@
In this paper we refer to such a cell complex as the prism complex of $\Gamma\mathcal{G}$ for convenience, however they are in fact a subclass of what Genevois calls prism complexes of quasi-median graphs [17]. In fact, what we call the prism complex is the 2–skeleton of Genevois’ quasi-median complex, which is obtained by filling every clique with a simplex, and filling every 1–skeleton of an nn–cube with an nn–cube. As we do not need to consider any higher dimensional cells, we omit them here.
@@ 2.4.2 Hyperplanes and carriers
A hyperplane may also be identified with its carrier: the union of closed cells that intersect the geometric hyperplane. From definitions one can verify that the 0–skeleton of a carrier is precisely some coset of $G_{\mathsf{Star}(v)}$, where $v$ is the vertex associated to the hyperplane [17, 16]. We also make use of its 1–skeleton, the full subgraph induced in $P(\Gamma\mathcal{G})$ by these vertices (following [38]). Note that the edges of the 1–skeleton that do not belong to the hyperplane itself are labeled by generators from $G_{\mathsf{Link}(v)}$.
@@ 2.4.3 Contact graph
The definition above is parallel to the construction of the contact graph of a CAT(0) cube complex, introduced by Hagen in [21]. Several results from the cube complex case also hold for graph products, in particular the following is known.
@@
[38, Corollary C] Let $\Gamma$ be a bounded degree simplicial graph, and let $\Gamma\mathcal{G}$ be a graph product of non-trivial groups. Then $\mathcal{C}P(\Gamma\mathcal{G})$ is a quasi-tree (and therefore a hyperbolic space), and the action of $\Gamma\mathcal{G}$ on $\mathcal{C}P(\Gamma\mathcal{G})$ is acylindrical.